if g(x)=7x-a where a is a constant and g(10)=3g(20), then g(30)=
Answer by Boreal(15235)

g(10)=70-a
g(20)=140-a so 3 g(20)=420-3a
but g(10)=3g(20) so 70-a=420-3a
2a=350
a=175
g(10)=70-175=-105
g(20)=140-175=-35
g(30)=210-175=35 ANSWER